a bacteria culture is known to grow at a rate proportional to the amount present. after one hour, 1000 strands of the bacteria are observed in the culture; and after four hours, 30000 strands. Find an expression fo the approximate number of strands of the bacteria present in the culture at any time t.


1
Expert's answer
2021-12-17T13:29:55-0500

Let "P(t)" be the size of the population of the bacteria culture at any time "t."

Then

"\dfrac{dP}{dt}=kP"

"\dfrac{dP}{P}=kdt"

"P(t)=P(0)e^{kt}"

Given "P(1)=1000, P(4)=30000"


"\dfrac{P(4)}{P(1)}=\dfrac{P(0)e^{4k}}{P(0)e^{k}}=\dfrac{30000}{1000}"

"e^{3k}=30"

"k=\dfrac{\ln 30}{3}"

"P(t)=P(0)e^{(\ln 30/3)t}"

"P(t)=P(0)(30)^{t\/3}"

"P(1)=P(0)(30)^{1\/3}=1000"

"P(0)=\dfrac{1000}{\sqrt[3]{30}}"

"P(t)=\dfrac{1000}{\sqrt[3]{30}}(30)^{t/3}"

"P(t)\approx322(30)^{t/3}, t\geq 0"